OpenStudy (anonymous):

what is the difference of the whole number and the mixed number below? a.4 2/5 b. 5 2/5 c.5 1/5 d. 4 3/5

OpenStudy (anonymous):

We need the whole and mixed number so we can find the difference...

OpenStudy (turingtest):

I sense... missing info. Or at least an unclear question

OpenStudy (anonymous):

oh im sorry 8 - 3 2/5

OpenStudy (turingtest):

8- 3 2/5=40/5-17/5=23/5

OpenStudy (turingtest):

=4 3/5 d.

OpenStudy (anonymous):

more info